Calculate the quantization noise for a digital
signal with step size 10 mV.
1
Expert's answer
2019-06-04T12:05:46-0400

Quantization noise is caused by rounding errors between the analog input voltage and nearest quantization level. Quantization noise power is related to step size as

"P = \frac {\Delta^2}{12}."

For step size

"\Delta = 10mV = 0.01\ V"

the noise power is equal to


"P = \frac {0.01^2}{12} = 8.33*10^{-6}\ W."